# SOPHY: Learning to Generate Simulation-Ready Objects with PHYsical Materials

### v2

@@ 1 Introduction
Generating high-quality 3D assets that can be incorporated in interactive virtual worlds or manufactured into functional 3D objects is a fundamental challenge in generative AI and digital content creation. Unfortunately, current generative models of 3D shapes [13, 89, 61] are limited to generating only static geometry and textures. Although there is significant effort in building 4D generative models for both geometry and motion synthesis [78, 66, 38], current methods are limited to fixed animations and are incapable of producing motions resulting from complex, dynamic object interactions.
In this work, we introduce a diffusion-based generative model for physics-aware 3D object synthesis, capable of generating 3D assets with detailed shape, texture, and, most importantly, physical material attributes governing kinematic object deformations (e.g., elastic deformations, plastic softening) and frictional interactions. Our approach produces physics-aware, interactive 3D objects, which are helpful for downstream applications such as physically based simulation, robotic interaction, and manufacturing. There are various challenges to developing such a generative approach. First, current 3D and 4D datasets do not contain physical material attributes for objects, making it hard to train, or even fine-tune such physics-aware generative models. Second, it is unclear how to jointly model the interplay of shape and material in 3D asset generation. Clearly, material attributes are not independent of geometric shapes, as repeatedly observed in prior works of physical material recognition from shapes  [37, 40, 63, 1]. For example, a chair’s backrest made of thin rods or strips is usually metal, while a thicker, block-like backrest is often made of fabric or soft padding.
@@ 2 Related Work
Over the recent years, we have witnessed an explosion of 3D generative models capable of generating detailed geometry and texture. The advances have been driven by expressive neural representations, such as NeRFs  [47, 54], Gaussian splats [27], signed distance fields [14, 52], occupancy fields [46], and numerous other representations for capturing shape or/and texture [49, 59, 21, 39, 82, 10, 11, 25, 85, 53, 18, 20, 81, 2] to name a few. We also refer readers to the recent surveys of 3D generative model [34, 75]. Unfortunately, the vast majority of existing 3D generative models neglect the physical material properties of the generated 3D objects, limiting their applicability to interactive and simulation-based environments. To address this gap, recent research has begun incorporating physical priors and constraints into the generative pipeline. For example, DiffuseBot [74] evaluates generated 3D robot models based on their simulation performance and refines the sampling distribution to favor more successful designs. Atlas3D [12] and Phys-Comp [19] enforce static equilibrium constraints during shape optimization, ensuring that generated objects remain stable under gravity. Despite these advances, all these prior works assume uniform or limited material properties, restricting their ability to generate diverse and physically intricate 3D assets. In contrast, our approach jointly optimizes 3D shape, albedo color, and material properties with a network that captures their natural dependencies.
@@
The advances in 3D generative models have also been led by the development of influential 3D datasets such as ShapeNet [7] and Objaverse [16]. More recent datasets, including ABO [15], Matsynth [71], and BlenderVault [42], incorporate surface texture information to train models for 3D generation and more plausible, photorealistic rendering. Despite this progress, a critical gap remains in the availability of detailed physical material information essential for accurately modeling material properties for physics-based simulations. While datasets such as ShapeNet-Mat [40], 3DCoMPaT [37] and its extensions [63, 1] attempt to address this limitation by providing part-level material labels, their annotations are coarse and unsuitable for direct use in physical simulators. To bridge this gap, we introduce a new dataset consisting of 3K objects spanning 12 shape categories, each annotated with precise part-level physical material properties.
@@
4D generation aims to create dynamic 3D content that aligns with input conditions such as images, text, or videos. Existing approaches primarily follow a data-driven pipeline [32, 76, 78, 38, 33, 65], leveraging off-the-shelf image or video diffusion models to generate dynamic scenes. For example, DreamGaussian4D [57] adopts Stable Video Diffusion [5] to animate 3D Gaussian splats (3DGS) [27] reconstructed from a single image. STAG4D [84] initializes multi-view images using image-to-image diffusion [60] anchored on input video frames from a text-to-video module. However, these methods rely on pre-trained generative models that lack physical understanding, often resulting in physically unrealistic motions [3]. To address this, PhysGaussian [80] and PhysDreamer [87] integrate the Material Point Method (MPM) [24] with 3DGS to produce physics-aware dynamics. Building on these efforts, several recent studies [41, 67, 9] have focused on physical dynamics generation from a single image. They typically follow a two-stage pipeline to obtain simulation-ready objects: generating 3D shapes first and then assigning them with homogeneous physical attributes by manual selection or estimation from VLMs. Despite offering an intuitive framework, they overlook the interdependence between geometries and material properties as they split shape generation and material estimation into two steps. Consequently, the assigned attributes may not accurately reflect the material heterogeneity and the tightly coupled nature of geometry observed in the real world. In contrast, our method introduces a learning-based approach for simulation-ready object generation. By jointly modeling shape geometries and material properties in an end-to-end framework, our method ensures greater physical realism and geometric coherence.
@@ 3 Material Annotation
We first describe the procedure for creating a dataset of 3D objects annotated with detailed physical material properties. Specifically, we annotate shapes with the material parameters used in the Material Point Method (MPM) [64, 24, 22], a popular simulator known for its effectiveness in handling complex simulations of behaviors and inertia effects inherent in solids  [80, 35, 6, 87], including but not limited to elastic and plastic deformation of solids and frictional interactions. The material parameters include: (a) Young’s modulus, (b) Poisson’s ratio, (c) yield stress, (d) friction angle, (e) density, and (f) material behavior type (also known as material model), including pure elastic deformation [77], plastic deformation with softening (damage) [72], plastic deformation without softening [73], and granular deformation [29].
Unfortunately, no datasets of 3D objects exist with such detailed material properties available. The most related dataset is 3DCoMPaT [37] along with its subsequent versions (3DCoMPaT++ [63], 3DCoMPaT200 [1]), which provide a rather coarse labeling of materials for 3D shape parts, such as “plastic”, “fabric”, “metal”, etc. While these categories provide some insight into the general physical characteristics of objects, they are often too broad for accurate simulation in physics engines. For example, a “plastic” part made out of rigid PVC is much stiffer (i.e., has a much higher Young’s modulus) than a flexible plastic material (e.g., LDPE). Unfortunately, such properties can be provided only by domain experts or manufacturers, and are difficult to interpret or estimate for annotators without a strong physics background and experience on commonly used material types in objects. Hiring experts to label these properties for every single component of every 3D object in a large database would be extremely cumbersome. Thus, we devised a semi-automatic pipeline that combines material property annotation by VLMs, followed by iterative expert feedback and verification.
@@ 3.1 VLM-guided Material Property Proposals
VLMs have recently become prominent in physical reasoning [90, 8, 30, 88, 41, 43] due to their extensive knowledge bases built on multi-modal data. Thus, we leveraged their zero-shot reasoning ability to give an initial estimation of material properties for input 3D shapes. We started by choosing $12$ shape categories from 3DCoMPaT200 [1], which contained a variety of material compositions and behaviors capable of elastic or plastic deformation, such as bags, pillows, and chairs. We skipped categories where objects behave rigidly (e.g., cabinets, faucets). As shown in Figure 2(top), we provided a popular VLM (ChatGPT-4o [50]) with (a) two automatically generated, rendered views of each textured object in the selected categories, (b) a textual description of the shape, including its object category, part tag, coarse material label, and color, as provided by 3DCoMPaT200, (c) a list of available fine-grained material categories e.g., for “plastic”, we include sub-types such as polypropylene, rigid or flexible PVC, and so on. We provide this fine-grained list in Section D.1. Our text prompt asks the VLM to identify the most plausible fine-grained material category per part. Then for each part, in a second round, we further prompted the VLM to provide the most likely material models and parameters (e.g., Young’s modulus, Poisson’s ratio, yield stress, and friction angle) based on a similar textual description, including this time the fine-grained material category as shown in Figure 2(bottom). This two-round prompt provided better material property prediction based on our expert verification.
@@ 3.2 Expert Verification and Feedback
We created five test scenarios to simulate object dynamics: (1) Dropping an object from a certain height; (2) Throwing an object in a certain direction; (3) Tilting an object; (4) Dragging an object; (5) Applying a short-term, time-variant (e.g., wind) force. We obtain object motions under these scenarios using a particle-based simulator [91] that we extended to model heterogeneous materials per particle, since our dataset often contains objects whose components are made of different materials. Then, we render object sequences with a high-performance renderer [31, 27] and obtain the simulation videos. After collecting these videos, we enlisted five mechanical engineering graduates to evaluate the physical plausibility of each video and instructed them to reach a consensus on their judgment. As shown in Figure 3(2)-(3a), for simulated objects deemed as realistic by this group, we store their material properties in our dataset. For objects deemed to have unrealistic motion, we ask the group to provide feedback to the VLM, specifying which part of the object seems to move in an implausible manner and why. This allows us to re-query VLMs for material parameters, as demonstrated in Figure 3(3b.1)-(3b.5), and generate new simulation videos automatically based on the updated parameters. This process is iterative until satisfactory simulation results are achieved.
@@ 3.3 Dataset Summary
After the above verification stage, we obtained $3,004$ 3D models in $12$ object categories originating from the 3DCoMPaT200 [1] dataset. Every shape part is labeled with physical material properties and detailed material categories, resulting in a total of $15,575$ labeled parts. We split our dataset into training, validation, and test sets, following the 3DCoMPaT200 splits – the total number of samples in each split is $2462$, $180$, and $362$, respectively. We provide more information about our dataset, including data statistics and sample simulation sequences in Appendix C.
@@ 4 Generative Model
Our proposed generative model, SOPHY, is based on Stable Diffusion [58], which generates data by denoising a compressed feature space (latent space). To extract this latent space, we process our shapes along with the color and physical material parameters through a variational autoencoder (Section 4.1). Then, a latent diffusion model is trained to jointly model the distribution of object geometries, colors, and material properties in this latent space (Section 4.2). The rationale behind jointly modeling these properties together is that they are strongly interdependent on each other i.e., certain shapes are strongly correlated with the use of specific materials and colors e.g., thin cantilever-shaped chair bases are associated with the use of metal and have a metallic gray color appearance. Modeling these properties independently of each other would yield unlikely or impossible materials for sampled geometries. After training our diffusion model, it can be sampled to generate new latent codes, which are then decoded to obtain shapes, compatible colors and physical material properties.
@@ 4.1 Shape, Texture and Material Autoencoder
Our autoencoder (Figure 4) compresses input 3D objects into compact latent codes that encode shape, color (albedo texture) and physical material properties. We discuss it below in more detail.
@@
[…]
@@
To aggregate the per-point information from the augmented point cloud, we design a set-to-set network inspired by 3DShape2VecSet [85] as our encoder. We first sub-sample a smaller point cloud $\boldsymbol{P}^{\prime}$ with $M=512$ fewer points from the original point cloud through furthest point sampling (FPS) [55], and augment it with color and material information to obtain $\boldsymbol{X}^{\prime}=\{\boldsymbol{x}^{\prime}_{i}\}_{i=1}^{M}$. Then, we use cross attention to produce the object’s latent representation:
@@
The above representation is quite high-dimensional. We thus seek to compress it towards a more compact latent code. Following Stable Diffusion [58] and 3DShape2VecSet [85], we adopt a variational autoencoder (VAE) regularized with the KL-divergence to achieve this effect. We first use two fully connected (FC) layers to project each latent code $\boldsymbol{f}_{i}$ to mean and variance:
@@
One possible decoder design follows the approach of 3DShape2VecSet [85], where latent codes are first processed through a series of self-attention blocks, then for a query point $\boldsymbol{q}\in\mathbb{R}^{3}$, its occupancy is determined by interpolating the latent codes based on the query position and transforming the result through a fully connected layer. This design could be naturally extended to also decode color and material information per query point. However, we observed worse performance using this approach, as color and material information are not meaningful for non-occupied query points (i.e., points outside the shape volume).
Empirically, we achieved better results with an alternative design. We first decode geometry in terms of occupancy values, then decode color, and finally decode material properties, but only for query points classified as occupied. Performance was further improved when we explicitly split the latent codes $\boldsymbol{z}_{i}$ into three sub-codes: the shape code $\boldsymbol{z}_{i,s}$ for occupancy decoding, the color code $\boldsymbol{z}_{i,c}$ for texture decoding, and the material code $\boldsymbol{z}_{i,m}$ for material property decoding. Each sub-code has a shape of $M\times\frac{C_{0}}{3}$. To decode these components, we introduce three dedicated branches, each responsible for processing one of the sub-codes (see Figure 4). We also apply cross-attention layers to the color and material branches, enabling them to attend to information from previously decoded latents. This design models the natural dependency of color on geometry and material properties on both geometry and color, which is inspired by the workflow commonly used in 3D asset creation, where artists typically start by defining the object’s shape, then apply textures, and finally assign physical attributes. Specifically, the cross-attention in our decoder is formulated as follows:
$\begin{gathered}\{\boldsymbol{f}^{(l)}_{i,s}\}=\mathrm{SelfAttn}\big{(}\{\boldsymbol{f}^{(l-1)}_{i,s}\}\big{)},\ \{\boldsymbol{f}^{(l)}_{i,c}\}=\mathrm{CrossAttn}\big{(}\{\boldsymbol{f}^{(l-1)}_{i,c}\},\{\boldsymbol{f}^{(l-1)}_{i,s}\}\big{)},\\ \{\boldsymbol{f}^{(l)}_{i,m}\}=\mathrm{CrossAttn}\big{(}\{\boldsymbol{f}^{(l-1)}_{i,m}\},\{\big{[}\boldsymbol{f}^{(l-1)}_{i,s},\boldsymbol{f}^{(l-1)}_{i,c}\big{]}\}\big{)},\end{gathered}$ (4)
where $l$ is the layer index, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation.
@@
where $\boldsymbol{f}_{\boldsymbol{q}}$ is a feature vector obtained by processing the query point through our encoder described in Equation 1. The functions $q(\cdot),k(\cdot),v(\cdot)$ are the query-key-value linear transformations used in attention [70]. By disentangling the latent codes, we ensure that occupancy is determined solely based on the shape sub-codes relevant to this task.
Color and material properties are decoded similarly, each using its own FC block and interpolation over the corresponding color and material sub-codes. For color prediction, we apply normalization after the FC block to ensure the output falls within the $[0,1]$ range. Young’s modulus and yield stress are always positive, so the FC blocks predict their values on a logarithmic scale. The friction angle has a range of $[0,\pi/2]$, so we apply a sigmoid activation after the FC block and scale the output by $\pi/2$. The categorical material behavior type is predicted using a softmax activation after the FC block. We store the material properties only for query points predicted as occupied. For color, we only predict the values for query points sampled on the mesh surface obtained by marching cubes [44].
@@
While training our variational autoencoder, we jointly optimize a combination of loss functions involving 3D occupancies, colors, and material properties. Specifically, we minimize a weighted sum of the following losses: (a) occupancy loss ${\mathcal{L}}_{o}$ (binary cross-entropy), (b) color loss ${\mathcal{L}}_{c}$ ($\ell_{1}$ norm), (c) Young’s modulus loss ${\mathcal{L}}_{E}$ ($\ell_{1}$ norm applied to logarithmic values), (d) Poisson’s ratio loss ${\mathcal{L}}_{\nu}$ ($\ell_{1}$ norm), (e) yield stress loss ${\mathcal{L}}_{\sigma}$ ($\ell_{1}$ norm applied to logarithmic values), (f) friction angle loss ${\mathcal{L}}_{\phi}$ ($\ell_{1}$ norm), (g) density loss ${\mathcal{L}}_{\rho}$ ($\ell_{1}$ norm), (h) material model loss ${\mathcal{L}}_{M}$ (cross-entropy), (i) a regularization loss ${\mathcal{L}}_{r}$ imposing a KL penalty towards a standard normal distribution on the latent codes, as typically used in VAEs [28]. We note that the color and material losses are computed only for training query points on or inside the object’s surface. For off-surface training points, we assign color and material properties by copying them from their nearest surface points. The combined loss function is:
@@
where $\lambda_{\omega}$ are weight parameters for balancing different terms. To further enhance training, we leverage the pretrained 3DShape2VecSet model [85], which was trained on ShapeNet [7] (a larger dataset containing 55K shapes) using occupancy supervision alone. Specifically, we initialize all network layers shared with 3DShape2VecSet, including the cross-attention weights on point positions in our encoder, the self-attention weights on shape sub-codes, and the occupancy decoder weights, using their pretrained values. This provides a small boost compared to training our model from scratch on our smaller 3K-shape dataset. We also note that we adopt the latent set representation of 3DShape2VecSet [85] in our autoencoder due to its simplicity and efficiency in compressing 3D objects. Our main insight is the joint modeling of shape, texture, and physical properties, and this concept is applicable beyond this specific representation.

### v3

@@ 1 Introduction
Generating high-quality 3D assets that can be incorporated in interactive virtual worlds or manufactured into functional 3D objects is a fundamental challenge in generative AI and digital content creation. Unfortunately, current generative models of 3D shapes chen20243dtopia ; zhao2025hunyuan3d ; siddiqui2024meta are limited to generating only static geometry and textures. Although there is significant effort in building 4D generative models for both geometry and motion synthesis wu2024cat4d ; sun2024dimensionx ; liangdiffusion4d , current methods are limited to fixed animations and are incapable of producing motions resulting from complex, dynamic object interactions.
In this work, we introduce a diffusion-based generative model for physics-aware 3D object synthesis, capable of generating 3D assets with detailed shape, texture, and, most importantly, physical material attributes governing kinematic object deformations (e.g., elastic deformations, plastic softening) and frictional interactions. Our approach produces physics-aware, interactive 3D objects, which are helpful for downstream applications such as physically based simulation, robotic interaction, and manufacturing. There are various challenges to developing such a generative approach. First, current 3D and 4D datasets do not contain physical material attributes for objects, making it hard to train, or even fine-tune such physics-aware generative models. Second, it is unclear how to jointly model the interplay of shape and material in 3D asset generation. Clearly, material attributes are not independent of geometric shapes, as repeatedly observed in prior works of physical material recognition from shapes  li20223d ; lin2018learning ; slim20233dcompat ; ahmed20253dcompat200 . For example, a chair’s backrest made of thin rods or strips is usually metal, while a thicker, block-like backrest is often made of fabric or soft padding.
@@ 2 Related Work
Over the recent years, we have witnessed an explosion of 3D generative models capable of generating detailed geometry and texture. The advances have been driven by expressive neural representations, such as NeRFs  mildenhall2020nerf ; Poole:2022:Dreamfusion , Gaussian splats kerbl20233d , signed distance fields chen2018implicit_decoder ; Park_2019_CVPR , occupancy fields Mescheder19 , and numerous other representations for capturing shape or/and texture oechsle2019texture ; shen2021deep ; hui2022wavelet ; Lin:2023:Magic3D ; xu2024grm ; chen2024meshanything ; chen2024meshanythingv2 ; jun2023shapegeneratingconditional3d ; zhang20233dshape2vecset ; petrov2024gem3d ; georgiou2025im2surftex ; hong2024lrm ; xu2024instantmesh ; bala2024edify to name a few. We also refer readers to the recent surveys of 3D generative model li2024advances3dgenerationsurvey ; wang2024diffusionmodels3dvision . Unfortunately, the vast majority of existing 3D generative models neglect the physical material properties of the generated 3D objects, limiting their applicability to interactive and simulation-based environments. To address this gap, recent research has begun incorporating physical priors and constraints into the generative pipeline. For example, DiffuseBot wang2023diffusebot evaluates generated 3D robot models based on their simulation performance and refines the sampling distribution to favor more successful designs. Atlas3D chen2024atlas3d and Phys-Comp guo2024physically enforce static equilibrium constraints during shape optimization, ensuring that generated objects remain stable under gravity. Despite these advances, all these prior works assume uniform or limited material properties, restricting their ability to generate diverse and physically intricate 3D assets. In contrast, our approach jointly optimizes 3D shape, albedo color, and material properties with a network that captures their natural dependencies.
@@
The advances in 3D generative models have also been led by the development of influential 3D datasets such as ShapeNet chang2015shapenet and Objaverse deitke2023objaverse . More recent datasets, including ABO collins2022abo , Matsynth vecchio2024matsynth , and BlenderVault litman2025materialfusion , incorporate surface texture information to train models for 3D generation and more plausible, photorealistic rendering. Despite this progress, a critical gap remains in the availability of detailed physical material information essential for accurately modeling material properties for physics-based simulations. While datasets such as ShapeNet-Mat lin2018learning , 3DCoMPaT li20223d and its extensions slim20233dcompat ; ahmed20253dcompat200 attempt to address this limitation by providing part-level material labels, their annotations are coarse and unsuitable for direct use in physical simulators. To bridge this gap, we introduce a new dataset consisting of 3K objects spanning 12 shape categories, each annotated with precise part-level physical material properties.
@@
4D generation aims to create dynamic 3D content that aligns with input conditions such as images, text, or videos. Existing approaches primarily follow a data-driven pipeline lee2024vividdream ; watson2024controlling ; wu2024cat4d ; liangdiffusion4d ; li20244k4dgen ; sun2025egd , leveraging off-the-shelf image or video diffusion models to generate dynamic scenes. For example, DreamGaussian4D ren2023dreamgaussian4d adopts Stable Video Diffusion blattmann2023stable to animate 3D Gaussian splats (3DGS) kerbl20233d reconstructed from a single image. STAG4D zeng2024stag4d initializes multi-view images using image-to-image diffusion shi2023zero123++ anchored on input video frames from a text-to-video module. However, these methods rely on pre-trained generative models that lack physical understanding, often resulting in physically unrealistic motions bansal2024videophy . Another line of research gao2025partrm ; liu2025riganything ; sun2025armo focuses on part-level motion based on articulation, yet is able to accommodate only rigid motion of parts. To address more general physics-aware dynamics, PhysGaussian xie2023physgaussian and PhysDreamer PhysDreamer integrate the Material Point Method (MPM) jiang2016material to animate 3DGS representations. Building on these efforts, several recent studies lin2024phys4dgen ; tan2024physmotion ; chen2025physgen3d ; liu2025unleashing have focused on physical dynamics generation from a single image. They typically follow a two-stage pipeline to obtain simulation-ready objects: generating 3D shapes first and then assigning them with homogeneous physical attributes by manual selection or estimation from VLMs. Despite offering an intuitive framework, they overlook the interdependence between geometries and material properties as they split shape generation and material estimation into two steps. Consequently, the assigned attributes may not accurately reflect the material heterogeneity and the tightly coupled nature of geometry observed in the real world. In contrast, our method introduces a learning-based approach for simulation-ready object generation. By jointly modeling shape geometries and material properties in an end-to-end framework, our method ensures greater physical realism and geometric coherence.
@@ 3 Material Annotation
We first describe the procedure for creating a dataset of 3D objects annotated with detailed physical material properties. Specifically, we annotate shapes with the material parameters used in the Material Point Method (MPM) sulsky1994particle ; jiang2016material ; hu2018moving , a popular simulator known for its effectiveness in handling complex simulations of behaviors and inertia effects inherent in solids  xie2023physgaussian ; pacnerf ; cao2024neuma ; PhysDreamer , including but not limited to elastic and plastic deformation of solids and frictional interactions. The material parameters include: (a) Young’s modulus, (b) Poisson’s ratio, (c) yield stress, (d) friction angle, (e) density, and (f) material behavior type (also known as material model), including pure elastic deformation wineman2005some , plastic deformation with softening (damage) wang2020material , plastic deformation without softening wang2019simulation , and granular deformation klar2016drucker .
Unfortunately, no datasets of 3D objects exist with such detailed material properties available. The most related dataset is 3DCoMPaT li20223d along with its subsequent versions (3DCoMPaT++ slim20233dcompat , 3DCoMPaT200 ahmed20253dcompat200 ), which provide a rather coarse labeling of materials for 3D shape parts, such as “plastic”, “fabric”, “metal”, etc. While these categories provide some insight into the general physical characteristics of objects, they are often too broad for accurate simulation in physics engines. For example, a “plastic” part made out of rigid PVC is much stiffer (i.e., has a much higher Young’s modulus) than a flexible plastic material (e.g., LDPE). Unfortunately, such properties can be provided only by domain experts or manufacturers, and are difficult to interpret or estimate for annotators without a strong physics background and experience on commonly used material types in objects. Hiring experts to label these properties for every single component of every 3D object in a large database would be extremely cumbersome. Thus, we devised a semi-automatic pipeline that combines material property annotation by VLMs, followed by iterative expert feedback and verification.
@@ 3.1 VLM-guided Material Property Proposals
VLMs have recently become prominent in physical reasoning zhen20243dvla ; chen2024commonsense ; lai2024vision ; zhao2024automated ; lin2024phys4dgen ; liu2024physgen due to their extensive knowledge bases built on multi-modal data. Thus, we leveraged their zero-shot reasoning ability to give an initial estimation of material properties for input 3D shapes. We started by choosing $12$ shape categories from 3DCoMPaT200 ahmed20253dcompat200 , which contained a variety of material compositions and behaviors capable of elastic or plastic deformation, such as bags, pillows, and chairs. We skipped categories where objects behave rigidly (e.g., cabinets, faucets). As shown in Figure˜2(top), we provided a popular VLM (ChatGPT-4o hurst2024gpt4o ) with (a) two automatically generated, rendered views of each textured object in the selected categories, (b) a textual description of the shape, including its object category, part tag, coarse material label, and color, as provided by 3DCoMPaT200, (c) a list of available fine-grained material categories e.g., for “plastic”, we include sub-types such as polypropylene, rigid or flexible PVC, and so on. We provide this fine-grained list in Section˜D.1. Our text prompt asks the VLM to identify the most plausible fine-grained material category per part. Then for each part, in a second round, we further prompted the VLM to provide the most likely material models and parameters (e.g., Young’s modulus, Poisson’s ratio, yield stress, and friction angle) based on a similar textual description, including this time the fine-grained material category as shown in Figure˜2(bottom). This two-round prompt provided better material property prediction based on our expert verification.
@@ 3.2 Expert Verification and Feedback
We created five test scenarios to simulate object dynamics: (1) Dropping an object from a certain height; (2) Throwing an object in a certain direction; (3) Tilting an object; (4) Dragging an object; (5) Applying a short-term, time-variant (e.g., wind) force. We obtain object motions under these scenarios using a particle-based simulator zong2023neural that we extended to model heterogeneous materials per particle, since our dataset often contains objects whose components are made of different materials. Then, we render object sequences with a high-performance renderer Laine2020diffrast ; kerbl20233d and obtain the simulation videos. After collecting these videos, we enlisted five mechanical engineering graduates to evaluate the physical plausibility of each video and instructed them to reach a consensus on their judgment. As shown in Figure˜3(2)-(3a), for simulated objects deemed as realistic by this group, we store their material properties in our dataset. For objects deemed to have unrealistic motion, we ask the group to provide feedback to the VLM, specifying which part of the object seems to move in an implausible manner and why. This allows us to re-query VLMs for material parameters, as demonstrated in Figure˜3(3b.1)-(3b.5), and generate new simulation videos automatically based on the updated parameters. This process is iterative until satisfactory simulation results are achieved.
@@ 3.3 Dataset Summary
After the above verification stage, we obtained $3,004$ 3D models in $12$ object categories originating from the 3DCoMPaT200 ahmed20253dcompat200 dataset. Every shape part is labeled with physical material properties and detailed material categories, resulting in a total of $15,575$ labeled parts. We split our dataset into training, validation, and test sets, following the 3DCoMPaT200 splits – the total number of samples in each split is $2462$, $180$, and $362$, respectively. We provide more information about our dataset, including data statistics and sample simulation sequences in Appendix˜C.
@@ 4 Generative Model
Our proposed generative model, SOPHY, is based on Stable Diffusion rombach2022high , which generates data by denoising a compressed feature space (latent space). To extract this latent space, we process our shapes along with the color and physical material parameters through a variational autoencoder (Section 4.1). Then, a latent diffusion model is trained to jointly model the distribution of object geometries, colors, and material properties in this latent space (Section 4.2). The rationale behind jointly modeling these properties together is that they are strongly interdependent on each other i.e., certain shapes are strongly correlated with the use of specific materials and colors e.g., thin cantilever-shaped chair bases are associated with the use of metal and have a metallic gray color appearance. Modeling these properties independently of each other would yield unlikely or impossible materials for sampled geometries. After training our diffusion model, it can be sampled to generate new latent codes, which are then decoded to obtain shapes, compatible colors and physical material properties.
@@ 4.1 Shape, Texture and Material Autoencoder
Our autoencoder (Figure˜4) compresses input 3D objects into compact latent codes that encode shape, color (albedo texture) and physical material properties. We discuss it below in more detail.
@@
We represent an object as a dense surface point cloud $\boldsymbol{P}=\{\boldsymbol{p}_{j}\}_{j=1}^{N}$, where $\boldsymbol{p}_{j}$ is a 3D point position ($N=2048$ in our implementation). The input to our autoencoder is a color- and material-augmented point cloud $\boldsymbol{X}=\{\boldsymbol{x}_{j}\}_{j=1}^{N}$, where each entry $\boldsymbol{x}_{j}$ contains the following per-point information: (a) 3D position $\boldsymbol{p}_{j}$, (b) RGB color $\boldsymbol{c}_{j}$, (c) a material property vector $\boldsymbol{m}_{j}$ with the following $9$ entries: (i) logarithm of Young’s modulus $E_{j}$ (a scalar), (ii) Poisson’s ratio $\nu_{j}$ (a scalar), (iii) logarithm of yield stress $\sigma_{j}$ (a scalar), (iv) friction angle $\phi_{j}$ (a scalar), (v) density $\rho_{j}$ (a scalar), and (vi) a $4$-dimensional learnable embedding $\mbox{{\boldmath$\mu$}}_{j}$ of material behavior type.
@@
To aggregate the per-point information from the augmented point cloud, we design a set-to-set network inspired by 3DShape2VecSet zhang20233dshape2vecset as our encoder. We first sub-sample a smaller point cloud $\boldsymbol{P}^{\prime}$ with $M=512$ fewer points from the original point cloud through furthest point sampling (FPS) pointnet++ , and augment it with color and material information to obtain $\boldsymbol{X}^{\prime}=\{\boldsymbol{x}^{\prime}_{i}\}_{i=1}^{M}$. Then, we use cross attention to produce the object’s latent representation:
@@
The above representation is quite high-dimensional. We thus seek to compress it towards a more compact latent code. Following Stable Diffusion rombach2022high and 3DShape2VecSet zhang20233dshape2vecset , we adopt a variational autoencoder (VAE) regularized with the KL-divergence to achieve this effect. We first use two fully connected (FC) layers to project each latent code $\boldsymbol{f}_{i}$ to mean and variance:
@@
One possible decoder design follows the approach of 3DShape2VecSet zhang20233dshape2vecset , where latent codes are first processed through a series of self-attention blocks, then for a query point $\boldsymbol{q}\in\mathbb{R}^{3}$, its occupancy is determined by interpolating the latent codes based on the query position and transforming the result through a fully connected layer. This design could be naturally extended to also decode color and material information per query point. However, we observed worse performance using this approach, as color and material information are not meaningful for non-occupied query points (i.e., points outside the shape volume).
Empirically, we achieved better results with an alternative design. We first decode geometry in terms of occupancy values, then decode color, and finally decode material properties, but only for query points classified as occupied. Performance was further improved when we explicitly split the latent codes $\boldsymbol{z}_{i}$ into three sub-codes: the shape code $\boldsymbol{z}_{i,s}$ for occupancy decoding, the color code $\boldsymbol{z}_{i,c}$ for texture decoding, and the material code $\boldsymbol{z}_{i,m}$ for material property decoding. Each sub-code has a shape of $M\times\frac{C_{0}}{3}$. To decode these components, we introduce three dedicated branches, each responsible for processing one of the sub-codes (see Figure˜4). We also apply cross-attention layers to the color and material branches, enabling them to attend to information from previously decoded latents. This design models the natural dependency of color on geometry and material properties on both geometry and color, which is inspired by the workflow commonly used in 3D asset creation, where artists typically start by defining the object’s shape, then apply textures, and finally assign physical attributes. Specifically, the cross-attention in our decoder is formulated as follows:
[…]
where $l$ is the layer index, and $\cdot$$\cdot$] denotes concatenation.
@@
where $\boldsymbol{f}_{\boldsymbol{q}}$ is a feature vector obtained by processing the query point through our encoder described in Equation˜1. The functions $q(\cdot),k(\cdot),v(\cdot)$ are the query-key-value linear transformations used in attention vaswani2017attention . By disentangling the latent codes, we ensure that occupancy is determined solely based on the shape sub-codes relevant to this task.
Color and material properties are decoded similarly, each using its own FC block and interpolation over the corresponding color and material sub-codes. For color prediction, we apply normalization after the FC block to ensure the output falls within the $[0,1]$ range. Young’s modulus and yield stress are always positive, so the FC blocks predict their values on a logarithmic scale. The friction angle has a range of $0,\pi/2]$, so we apply a sigmoid activation after the FC block and scale the output by $\pi/2$. The categorical material behavior type is predicted using a softmax activation after the FC block. We store the material properties only for query points predicted as occupied. For color, we only predict the values for query points sampled on the mesh surface obtained by marching cubes lorensen1987marching .
@@
While training our variational autoencoder, we jointly optimize a combination of loss functions involving 3D occupancies, colors, and material properties. Specifically, we minimize a weighted sum of the following losses: (a) occupancy loss ${\mathcal{L}}_{o}$ (binary cross-entropy), (b) color loss ${\mathcal{L}}_{c}$ ($\ell_{1}$ norm), (c) Young’s modulus loss ${\mathcal{L}}_{E}$ ($\ell_{1}$ norm applied to logarithmic values), (d) Poisson’s ratio loss ${\mathcal{L}}_{\nu}$ ($\ell_{1}$ norm), (e) yield stress loss ${\mathcal{L}}_{\sigma}$ ($\ell_{1}$ norm applied to logarithmic values), (f) friction angle loss ${\mathcal{L}}_{\phi}$ ($\ell_{1}$ norm), (g) density loss ${\mathcal{L}}_{\rho}$ ($\ell_{1}$ norm), (h) material model loss ${\mathcal{L}}_{M}$ (cross-entropy), (i) a regularization loss ${\mathcal{L}}_{r}$ imposing a KL penalty towards a standard normal distribution on the latent codes, as typically used in VAEs kingma2013auto . We note that the color and material losses are computed only for training query points on or inside the object’s surface. For off-surface training points, we assign color and material properties by copying them from their nearest surface points. The combined loss function is:
@@
where $\lambda_{\omega}$ are weight parameters for balancing different terms. To further enhance training, we leverage the pretrained 3DShape2VecSet model zhang20233dshape2vecset , which was trained on ShapeNet chang2015shapenet (a larger dataset containing 55K shapes) using occupancy supervision alone. Specifically, we initialize all network layers shared with 3DShape2VecSet, including the cross-attention weights on point positions in our encoder, the self-attention weights on shape sub-codes, and the occupancy decoder weights, using their pretrained values. This provides a small boost compared to training our model from scratch on our smaller 3K-shape dataset. We also note that we adopt the latent set representation of 3DShape2VecSet zhang20233dshape2vecset in our autoencoder due to its simplicity and efficiency in compressing 3D objects. Our main insight is the joint modeling of shape, texture, and physical properties, and this concept is applicable beyond this specific representation.